OpenStudy (anonymous):

How can I solve this ?? Find the values of x for which f(x)=g(x) 1.) f(x)=x2 , g(x)= x + 6 PLS.HELP ME :(((

myininaya (myininaya):

f(x)=g(x) x^2=x+6 x^2-x-6=0 (x-3)(x+2)=0 x=3 x=-2

OpenStudy (anonymous):

How did you get the x=3 and x= -2 ??

myininaya (myininaya):

if (x-3)(x+2)=0 then x-3=0 and/or x+2=0 x-3=0 implies x=3 x+2=0 implies x=-2

OpenStudy (anonymous):

Thank You :))

myininaya (myininaya):

you totally got it?

OpenStudy (anonymous):

Not really :)) Is x=3 and x=-2 are the answers ??